in standard form what is an equation that is perpendicular to 5x+y=4 and goes through points (8,9)?
Answer by mananth(16946)

5x+y = 4
Find the slope of this line

y=-5x+4
Compare this equation with y=mx+b
slope m = -5

The slope of a line perpendicular to the above line will be the negative reciprocal 1/5
m1*m2=-1


m= 1/5 ,point (8,9)
Find b by plugging the values of m & the point in
y=mx+b
9=8/5+b
b= 37/5
m= 1/5
The required equation is y= (1/ 5)x+37/ 5
m.ananth@hotmail.ca